OpenStudy (anonymous):

the numbers in this sequence increase by 7 each time. 1,8,15,22,29.... the sequence continues in the same way. will the number 777 be in this sequence? YES/NO Why?

OpenStudy (dayakar):

it is arithmetic progression

OpenStudy (theloshua):

no

OpenStudy (dayakar):

first term a=1 common difference =8-1 =7 d=7

OpenStudy (anonymous):

No because 1+111 X 7= 1=777=778 8+111 X 7= 8=777=785 785-7=778 It's one more than 777 So, the solution is NO. Hope it helps you.

OpenStudy (dayakar):

last term =777 (which is have to check)

OpenStudy (theloshua):

lol I already said no

OpenStudy (dayakar):

a+(n-1)d =777 1+(n-1)7 =777 he needs explanation

OpenStudy (anonymous):

No Any number in the series minus 1 should be divisible by 7,Like 29-1=28, 22-1=21 but 777-1=776 not divisible by 7!

OpenStudy (dayakar):

very good, best concept

OpenStudy (loser66):

The difference between numbers is 7, hence the general formula for it is \(a_n = 1+ 7(n-1) = 7n -6\) if 777 is in the sequence , then 777 = 7n -6, n = 771/7 = 110.1428 which is not integer. Hence 777 is not in the sequence.